Check whether the function f(x,y)={4x^2y/4x^4+y^2, (x,y) is not equal to (0,0)
0 , (x,y)=(0,0) is continuous at (0,0).
1
Expert's answer
2020-05-13T18:54:33-0400

Given

"f(x,y) = \frac{4x^2y }{4x^4+y^2}"

To evaluate

"\lim_{(x,y)\to(0,0)}f(x,y) = \lim_{(x,y)\to(0,0)}\frac{4x^2y }{4x^4+y^2}"

Choose path passes through (0,0), let

"y=mx"

then the limit along y=mx given by

"\begin{aligned}
 \lim_{(x,y)\to(0,0)}\frac{4x^2y }{4x^4+y^2}&= \lim_{x\to0}\frac{4mx^3 }{4x^4+m^2x^2}\\
&= \lim_{x\to0}\frac{4mx }{4x^2+m^2 }\\
&= \lim_{x\to0}\frac{0 }{0+m^2 }=0
\end{aligned}"

Since the limit does not depend on m , then limit exist and equal to 0

"\lim_{(x,y)\to(0,0)}f(x,y) =f(0,0)=0"

Then f(x,y ) is continuous at (0,0)